OpenStudy (anonymous):

plz convert 1AB hexadecimal to decimal.

OpenStudy (anonymous):

is it 256+10+11 =277

OpenStudy (anonymous):

does 1 correspond to 256?

OpenStudy (anonymous):

\[(1AB )_{16} = 1\times 16^2 + 10 \times 16^1 + 11\times 16^0 = (427)_{10} \]